OpenStudy (anonymous):

what is the volume of a cone that has a base with a diameter of 14 cm and a height of 8cm

OpenStudy (anonymous):

formula is \[\pi r^{2}h \div 3\]

OpenStudy (anonymous):

it wont show on the list thoe a:1230.88 b: 410.5 c:58.61 d:615.75

OpenStudy (anonymous):

I got one of the answers using the formula. All you have to do is plug them in Diameter= 2*r 14= 2*r 14/2=r 7=r and we know h=8 just plug that in into \[\frac{1}{3} (\pi*r^2*h)\]